Each year, my student’s Mike’s kazoo collection appreciates (goes up in value) 10%. Mike says that after 10 years, the value will double. How do you respond?


Answer by richwmiller(17219)

Let's take a simple example.
Let's say that this year the collection is worth $100
1.10*x equals the original x plus 10%
1- 1.1*100=110
2- 1.1*110=121
3- 1.1*121=133.10
4- 1.1*133.1=146.41
5- 1.1*146.41=161.051
6- 1.1*161.051=177.1561
7- 1.1*177.1561=194.872
8- 1.1*194.872=214.3592
After 8 years it has more than doubled.
100*1.10^8=214.359